# On the Inclusion of Charge and Spin States in Cartesian Tensor Neural Network Potentials

### v1

@@ Abstract
In this letter, we present an extension to TensorNet, a state-of-the-art equivariant Cartesian tensor neural network potential, allowing it to handle charged molecules and spin states without architectural changes or increased costs. By incorporating these attributes, we address input degeneracy issues, enhancing the model’s predictive accuracy across diverse chemical systems. This advancement significantly broadens TensorNet’s applicability, maintaining its efficiency and accuracy.
@@ 1 Introduction
Neural network potentials1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, which are deep learning models used to predict the energy and forces of atomic systems12, typically rely on atomic numbers and positions as their sole inputs. This simplification, while facilitating model construction and computation, overlooks other significant properties such as total charge and spin state, which are crucial attributes that influence the behavior and interaction of atomistic systems. Furthermore, neglecting total charge and spin state can lead to a degeneracy of inputs13. This degeneracy occurs when different molecular states, distinguishable only by their charge or spin, are represented by identical inputs for the model, impeding the model’s ability to discern the states and accurately predict potential energies and forces.
In the literature, charges have typically been addressed in neural network potentials with the aim of incorporating long-range electrostatics5, 14, 15, 16, 17. These approaches provide mechanisms for the global redistribution of charge across the system, utilizing neural networks informed by the chemical environment to predict intermediate properties, such as electronegativities and effective charges, and equilibration schemes18 that involve solving systems of linear equations 14, 15, 16, 13, 17, or through self-consistent processes19. Furthermore, in some cases, additionally predicted quantities such as atomic charges are used for the computation of manually introduced physics-based terms in the energy prediction5, 13.
In this work, we introduce a simple yet effective extension to TensorNet20, a state-of-the-art equivariant neural network potential based on Cartesian tensors, enabling it to accommodate charged molecules and spin states. This modification does not require any architectural changes to the original model, nor does it lead to increased computational costs. Even though the approach does not explicitly account for global redistribution of properties, this enhancement retains the model’s original predictive accuracy on the systems we studied, and the approach circumvents the need for explicit physical terms or the prediction of additional quantities. This enhancement expands TensorNet’s applicability to a broader range of chemical systems.
@@ 2 Method
One of the key operations in TensorNet is how neighboring nodes’ tensor features are aggregated and used together with the ones of the receiving node to generate a new set of node tensor representations that transform appropriately under $O(3)$. We refer the reader to Ref 20 for full details. In each layer, after some node level transformations $X^{(i)}\rightarrow X^{\prime(i)}=X^{(i)}/(||X^{(i)}||+1)\rightarrow Y^{(i)}$, pair-wise messages $M^{(ij)}$ from neighbor $(j)$ to receiving atom $(i)$ are built by decomposing neighbor’s features $Y^{(j)}$,
@@
where ${f_{I}}^{ij},{f_{A}}^{ij},{f_{S}}^{ij}$ are learnable functions of the distance between atoms $r_{ij}$, and $\phi(r_{ij})$ is a cosine cutoff function. Messages are summed for all neighbors, $M^{(i)}=\sum_{j}{M^{(ij)}}$, and the generation of new features $Y^{\prime(i)}$ from current node features $Y^{(i)}$ and message $M^{(i)}$ is performed via simple matrix product as
@@
effectively mixing scalars, vectors, and tensors, and ensuring $O(3)$-equivariance, as proved in Ref 20. Resulting representations are eventually manipulated to yield residual updates $\Delta X^{\prime(i)}$ to the layer’s input normalized features $X^{\prime(i)}$
@@
We propose to include molecular states’ information $\{\psi_{k}\}$, labeled by $k$, by modifying Equations (2) and (3) in TensorNet’s interaction layers in the following node-wise manner
$Y^{\prime(i)}=(1+\sum_{k}{\lambda_{k}\psi_{k}})(Y^{(i)}M^{(i)}+M^{(i)}Y^{(i)}),$ (4)
$X^{(i)}\leftarrow X^{\prime(i)}+\Delta X^{\prime(i)}+(1+\sum_{k}{\tilde{\lambda}_{k}\psi_{k}})(\Delta X^{\prime(i)})^{2}$ (5)
where ${\lambda}_{k},\tilde{\lambda}_{k}$ can be regarded as per-layer constant or learnable weights for the encoding of states such as total charge or spin $\psi_{k}=\{Q,S,...\}$. Notice that the modification does not break equivariance under $O(3)$, since it amounts to a rescaling of tensor features by means of a state-dependent scalar factor. Furthermore, it has been designed in such a way that when the system is neutral and in its singlet state, i.e.
@@
the model defaults to the original TensorNet (Eqs (2) and (3)), therefore recovering the predictive accuracy already demonstrated in Ref 20.
It is also worth mentioning that, since the inclusion of states is performed node-wise, one can also incorporate atomic attributes, such as partial charges ${q^{(i)}}$. In this case, one needs to rely on some external partial charge computation scheme, which might assume the availability of information beyond atomic numbers and positions, such as SMILES representations, molecular bonds’ topology or other preprocessing steps.
@@ 3 Results and discussion
We performed a series of experiments with the extended TensorNet model using the TorchMD-Net framework21. We trained the model on a custom-made dataset and on two publicly available datasets, targeting atomic number and geometry degeneracy, as well as the simultaneous presence of neutral and charged molecules regardless of structural or geometric overlap between these. We used a generic set of reasonable hyperparameters, without addressing dataset-dependent fine-tuning to obtain the best possible performances. In all cases, we used the direct model prediction of energies and forces (when needed), without the manual inclusion of physics-based terms. Overall, we evaluate TensorNet’s extension when dealing with total charges, Gasteiger partial charges, or singlet and triplet states.
@@ 3.1 Toy degeneracy problem
As a first test, to illustrate the input degeneracy issue and how TensorNet’s extension resolves it without affecting its baseline accuracy, we constructed two toy datasets, Dataset A and Dataset B, each one comprising five members of five pairs of unique molecular systems, each of the elements in the pair differing in total charge (see Fig 1). These pairs are indistinguishable to a neural network that does not account for total charge, as they share identical atomic numbers and geometric configurations. The datasets include total charges, Gasteiger partial charges computed with RDKit, and calculated energies and atomic forces for 2000 conformers per molecule using GFN2-xTB22. Therefore, each dataset contains a total of 10k data points. Conformers were generated by minimizing each molecule, displacing atomic positions with Gaussian noise with a standard deviation of 0.2 Å, and filtering them such that maximum atomic forces are <100 eV/Å.
@@
Splits of 50/10/40% were used for training, validation, and testing, respectively, both for single and combined datasets. For the charge-aware case, we tried separately the inclusion of total charges $Q$ and Gasteiger partial charges $q^{(i)}$, both with $\lambda_{Q}=\tilde{\lambda}_{Q}$ and $\lambda_{q^{(i)}}=\tilde{\lambda}_{{q^{(i)}}}$ equal to 0.1 across all layers. Results can be found in Table 1. As expected, the original model trained on the merged dataset has a very poor accuracy (TensorNet, Dataset A $\cup$ B). The extension with total and partial charges allows us to learn the merged dataset with sub-meV and sub-meV/Å differences in energy and force errors with respect to single dataset training of the corresponding extended model. Moreover, the use of total charge $Q$ gives better results than the use of partial charges $q^{(i)}$, since errors on the merged dataset and dataset A are lower.
Table 1: Toy datasets results. Energy (E) and forces (F) mean absolute errors in meV and meV/Å.
@@
Figure 1: Molecules included in the A and B toy datasets, from which 2,000 data points per molecule are obtained by generating conformers and computing potential energies and atomic forces using GFN2-xTB22. Columns illustrate degenerate pairs of molecules for a neural network that uses solely atomic numbers and positions as inputs.
@@ 3.2 SPICE PubChem
The PubChem subset inside the SPICE dataset23 (version 1.1.4) consists of conformers of small drug-like molecules, with energies and forces computed at the $\omega$B97M-D3(BJ)/def2-TZVPPD level of theory. Approximately 4% of the molecules and conformers in the dataset are not neutral. The dataset already provides total charges, while we computed Gasteiger partial charges for all the molecules in the subset, including the neutral ones. The computation of partial charges failed with RDKit for 42 molecules, with corresponding conformers being discarded. This resulted in 707,558 data points in total.
To include total charge $Q$ or atomic partial charges $q^{(i)}$, we take $\lambda_{Q}=\tilde{\lambda}_{Q}$ and $\lambda_{q^{(i)}}=\tilde{\lambda}_{{q^{(i)}}}$ constant and non-learnable across all layers, equal to 0.1 or 0.25. We used 80/10/10% splits for training, validation, and testing, respectively. In Table 2, we show that including total charge $Q$ or partial charges $q^{(i)}$ can improve energy and forces accuracy. Even though charged species comprise only 4% of the data points, the freedom given to the model to adapt to varying charge states improves significantly the accuracy on both charged and neutral conformations. With the best method, which uses total charge $Q$ with $\lambda_{Q}=\tilde{\lambda}_{Q}=0.1$, reductions of 30% and 12% in energy errors are achieved for charged and neutral test molecules, respectively. Nevertheless, the inclusion of Gasteiger partial charges gives relevant out-of-the-box improvements too.

### v2

# Broadening the Scope of Neural Network Potentials through Direct Inclusion of Additional Molecular Attributes

@@ Abstract
Most state-of-the-art neural network potentials do not account for molecular attributes other than atomic numbers and positions, which limits its range of applicability by design. In this work, we demonstrate the importance of including additional electronic attributes in neural network potential representations with a minimal architectural change to TensorNet, a state-of-the-art equivariant model based on Cartesian rank-2 tensor representations. By performing experiments on both custom-made and public benchmarking datasets, we show that this modification resolves the input degeneracy issues stemming from the use of atomic numbers and positions alone, while enhancing the model’s predictive accuracy across diverse chemical systems with different charge or spin states. This is accomplished without tailored strategies or the inclusion of physics-based energy terms, while maintaining efficiency and accuracy. These findings should furthermore encourage researchers to train and use models incorporating these additional representations.
@@ 1 Introduction
Neural network potentials1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, which are deep learning models used to predict the energy and forces of atomic systems12, have emerged as a tool making possible to circumvent the trade-off between computational cost and accuracy in computational chemistry. Despite its remarkable successes, these architectures typically rely on atomic numbers and positions as their only inputs, resulting in a simplification that overlooks significant quantum properties such as total charge and spin state by design. Neglecting these properties can lead to a degeneracy of inputs, where different molecular states, distinguishable only by their charge or spin, are represented identically by the neural network.
Given these considerations, the quest for extending the range of applicability of these models and for building general neural network potentials requires strategies making the incorporation of these additional attributes possible. For example, as a consequence of these current limitations, researchers have sometimes curated real-world datasets 13 to avoid the inclusion of charged species 14, thereby restricting the diversity and realism of the training data, or directly avoided a special treatment of these attributes15. To overcome these challenges, it is essential to integrate these properties 16, 17, 18.
The incorporation of charges has been previously addressed in the literature to take into account the long-range effects that they induce in molecular systems5, 19, 20, 21, 22, 23, 24, 18. Previous approaches have typically involved mechanisms for the global redistribution of charge across the system, with neural networks informed by the chemical environment that predict intermediate properties such as electronegativities and effective charges, and equilibration schemes25 that solve systems of linear equations19, 20, 21, 22 or use self-consistent processes26. In other models5, 16 additional quantities interpreted as atomic charges are predicted and used in ad-hoc introduced physics-based terms for energy prediction, such as Coulomb potential energies. Recently, the Latent Ewald Summation method 27, 28 proposed learning hidden charges from local features to compute long-range interactions, avoiding explicit charge equilibration while maintaining accuracy. While this approach can naturally handle long-range effects, an electrostatic-like interaction in reciprocal space is assumed.
In this work, we introduce a simple yet effective extension to a state-of-the-art model such as TensorNet29, allowing it to accommodate charged molecules and spin states without requiring architectural changes or additional featurizations, and therefore breaking the aforementioned degeneracy problem that most of the current models display. Despite not explicitly accounting for the global redistribution of properties, this enhancement retains the model’s original predictive accuracy without using additional physical terms or predicted quantities, being able to achieve state-of-the-art accuracy in common benchmark datasets. On top of this, since the presented method allows one to include these additional attributes in the learned representations, the prediction of other quantities used in the literature for the incorporation of long-range interactions is facilitated. Overall, the modification of TensorNet presented here expands its applicability to a broader range of chemical systems, addressing a critical shortcoming of most state-of-the-art architectures.
@@ 2 Method
One of the key operations in TensorNet, distinguishing it from other higher-rank spherical equivariant models, is how neighboring nodes’ tensor features are aggregated and used together with the ones of the receiving node to generate a new set of node tensor representations that transform appropriately under $O(3)$. We refer the reader to Ref 29 for full details and derivations. In each layer, after some node level transformations $X^{(i)}\rightarrow X^{\prime(i)}=X^{(i)}/(||X^{(i)}||+1)\rightarrow Y^{(i)}$, pair-wise messages $M^{(ij)}$ from neighbor $(j)$ to receiving atom $(i)$ are built by decomposing neighbor’s features $Y^{(j)}$,
@@
where ${f_{I}}^{ij},{f_{A}}^{ij},{f_{S}}^{ij}$ are learnable functions of the distance between atoms $r_{ij}$, and $\phi(r_{ij})$ is a cosine cutoff function. Tensor messages are summed for all neighbors,
$M^{(i)}=\sum_{j\in\mathcal{N}(i)}{M^{(ij)}}$ (2)
and the generation of new features $Y^{\prime(i)}$ from current node features $Y^{(i)}$ and corresponding message $M^{(i)}$ is performed via simple matrix product as
@@
effectively mixing scalars, vectors, and tensors, and ensuring $O(3)$-equivariance, as proved in Ref 29. Resulting representations are eventually manipulated to yield residual updates $\Delta X^{\prime(i)}$ to the layer’s input normalized features $X^{\prime(i)}$
@@
We propose to include molecular states’ information $\psi_{k}$, by modifying Equations (2) and (3) in TensorNet’s interaction layers in the following node-wise manner
$Y^{\prime(i)}=(1+{\lambda_{k}\psi_{k}})(Y^{(i)}M^{(i)}+M^{(i)}Y^{(i)}),$ (5)
$X^{(i)}\leftarrow X^{\prime(i)}+\Delta X^{\prime(i)}+(1+{\tilde{\lambda}_{k}\psi_{k}})(\Delta X^{\prime(i)})^{2}$ (6)
where ${\lambda}_{k},\tilde{\lambda}_{k}$ can be regarded as per-layer constant or learnable weights for the encoding of states such as total charge or spin $\psi_{k}=\{Q,S,...\}$. Notice that the modification, while forcing the network to output different values when considering the same input atomic numbers and positions, does not break equivariance under $O(3)$, since it amounts to a rescaling of tensor features by means of a state-dependent scalar factor. Furthermore, it has been designed in such a way that when the system is neutral and in its singlet state, i.e.
@@
the model defaults to the original TensorNet (Eqs (2) and (3)), therefore recovering the predictive accuracy already demonstrated in Ref 29.
Note that since operations are performed node-wise, this framework naturally accommodates both atomic and molecular attributes. For atomic attributes (such as partial charges $\psi_{k}={q^{(i)}}$), each node’s operations can be modified by its specific value. For molecular attributes (like total charge $\psi_{k}=Q$ or spin $\psi_{k}=S$), the same value modifies operations across all nodes in the system. This allows incorporating either local or global information within the same mathematical framework. In the case of atomic charges, commonly used in biomolecular simulations, one needs to rely on some external partial charge computation scheme, which might assume the availability of information beyond atomic numbers and positions, such as SMILES representations, molecular bonds or other preprocessing steps. Furthermore, given the possibility of including total charge as an input to the network, our method also enables the prediction of additional atomic quantities that can be trained to match DFT partial charges and used to evaluate electrostatic terms, as previously done in the literature16, 21, 22. Nevertheless, we won’t make use of this capability in our experiments, showing that highly accurate models such as TensorNet do not require the inclusion of physics-based terms, such as Coulomb, dispersion or nuclear repulsion, for the benchmark systems studied in this work and commonly used in the literature.
Note that while this framework theoretically supports multiple attributes, careful consideration would be needed to handle them simultaneously. For example, one could consider the modification of Equations 5 and 6, and promote $\lambda_{k}\psi_{k}$ to a linear combination of attributes $\sum_{k}\lambda_{k}\psi_{k}$. However, fixed weights could lead to degeneracies, where distinct quantum states produce identical scalar factors. A more sophisticated approach using learnable weighting parameters $\lambda_{k}$ could help discover encodings that properly distinguish different quantum states while preserving their physical influence. The current implementation demonstrates the effectiveness of individual attributes separately, leaving the exploration of simultaneous quantum attributes as future work.
@@ 3 Results and discussion
We performed a series of experiments with the extended TensorNet model using the TorchMD-Net framework30. We trained the model on a custom-made dataset and several publicly available datasets, targeting atomic number and geometry degeneracy, as well as the simultaneous presence of different charge states regardless of structural or geometric overlap between these. We used a generic set of reasonable hyperparameters, without addressing dataset-dependent fine-tuning to obtain the best possible performances. In all cases, we used the direct model prediction of energies and forces (when needed), without the inclusion of physics-based terms. Overall, we evaluate TensorNet’s extension when dealing with total charges, Gasteiger partial charges, or singlet and triplet states. Training details and hyperparameters used in the experiments can be found in the Supplementary Information.
@@ 3.1 Toy degeneracy problem
As a first test, to illustrate the input degeneracy issue and how TensorNet’s extension resolves it without affecting its baseline accuracy, we constructed two toy datasets, Dataset A and Dataset B, each one comprising five members of five pairs of unique molecular systems, each of the elements in the pair differing in total charge (see Fig 1). These pairs are indistinguishable from a neural network that does not account for total charge, as they share identical atomic numbers and geometric configurations. The datasets include total charges, Gasteiger partial charges computed with RDKit, and calculated energies and atomic forces for 2000 conformers per molecule using GFN2-xTB31. Therefore, each dataset contains a total of 10k data points. Conformers were generated by minimizing each molecule, displacing atomic positions with Gaussian noise with a standard deviation of 0.2 Å, and filtering them such that maximum atomic forces are <100 eV/Å.
@@
Splits of 50/10/40% were used for training, validation, and testing, respectively, both for single and combined datasets. For the charge-aware case, we tried separately the inclusion of total charges $Q$ and Gasteiger partial charges $q^{(i)}$, both with $\lambda_{Q}=\tilde{\lambda}_{Q}$ and $\lambda_{q^{(i)}}=\tilde{\lambda}_{{q^{(i)}}}$ equal to 0.1 across all layers. Results can be found in Table 1. As expected, the original model trained on the merged dataset has a very poor accuracy (TensorNet, Dataset A $\cup$ B). The extension with total and partial charges allows us to learn the merged dataset with sub-meV and sub-meV/Å differences in energy and force errors with respect to single dataset training of the corresponding extended model. The use of total charge $Q$ gives better results than the use of partial charges $q^{(i)}$.
Table 1: Toy datasets results. Energy (E) and forces (F) mean absolute errors in meV and meV/Å. $Q$ refers to total charge while $q^{(i)}$ refers to Gasteiger partial charges, with factors corresponding to weights $\lambda_{Q}=\tilde{\lambda}_{Q}$ and $\lambda_{q^{(i)}}=\tilde{\lambda}_{{q^{(i)}}}$, respectively.
@@
Figure 1: Molecules included in the A and B toy datasets, from which 2,000 data points per molecule are obtained by generating conformers and computing potential energies and atomic forces using GFN2-xTB31. Columns illustrate degenerate pairs of molecules for a neural network that uses solely atomic numbers and positions as inputs.
@@ 3.2 SPICE PubChem
The PubChem subset inside the SPICE dataset13 (version 1.1.4) consists of conformers of small drug-like molecules, with energies and forces computed at the $\omega$B97M-D3(BJ)/def2-TZVPPD level of theory. Approximately 4% of the molecules and conformers in the dataset are not neutral. The dataset already provides total charges, while we computed Gasteiger partial charges for all the molecules in the subset, including the neutral ones. The computation of partial charges failed with RDKit for 42 molecules, with corresponding conformers being discarded. This resulted in 707,558 data points in total. To include total charge $Q$ or atomic partial charges $q^{(i)}$, we take $\lambda_{Q}=\tilde{\lambda}_{Q}$ and $\lambda_{q^{(i)}}=\tilde{\lambda}_{{q^{(i)}}}$ constant and non-learnable across all layers, equal to 0.1 or 0.25. We used 80/10/10% splits for training, validation, and testing, respectively.
In Table 2, we show that including total charge $Q$ or partial charges $q^{(i)}$ can improve energy and forces accuracy. Even though charged species comprise only 4% of the data points, the freedom given to the model to adapt to varying charge states improves significantly the accuracy on both charged and neutral conformations. With the best method, which uses total charge $Q$ with $\lambda_{Q}=\tilde{\lambda}_{Q}=0.1$, reductions of 30% and 12% in energy errors are achieved for charged and neutral test molecules, respectively. The persisting performance gap between neutral and charged molecules (Table 2) can be attributed to dataset imbalance. This is evidenced by our toy dataset experiments (Table 1), where an equal representation of charge states (2000 conformations each) shows no accuracy degradation for the merged dataset. In fact, force prediction accuracy for the merged dataset falls between that of individual datasets, indicating comparable performance across charge states when properly balanced. The imbalanced case can be understood as model parameters being predominantly trained on $Q=0$ data, with a limited representation of other charge values. Nevertheless, even with just 4% charged species, including total charge consistently improves overall and charge-specific accuracy. This consistent improvement in prediction accuracy for both neutral and charged species suggests that the development of distinct internal representations due to charge awareness benefits the model beyond just handling charged molecules. This demonstrates that incorporating attributes like total charge helps the model develop more physically meaningful representations, enhancing predictive performance across all molecular states. The inclusion of Gasteiger partial charges gives relevant out-of-the-box improvements too.
@@
3.3 Carbon chain, silver clusters and sodium chloride clusters
These datasets were introduced in Ref 21 to illustrate the shortcomings of previous generations of models 32, 33, and are typically used as benchmarks for architectures incorporating long-range interactions and electrostatics, allowing us to compare their performances against the modified TensorNet, which does not incorporate physics-based terms in the energy prediction.
Carbon chain, $\mathbf{C_{10}H_{2}/C_{10}H_{3}^{+}}$. This benchmark consists of two subsystems: a neutral linear carbon chain with hydrogen atoms at each end, $\mathrm{C_{10}H_{2}}$, and a protonated counterpart, $\mathrm{C_{10}H_{3}^{+}}$, yielding a total charge of +1 and resulting in a global charge redistribution.
Silver clusters, $\mathbf{Ag_{3}^{+/-}}$. In this case, the dataset consists of triangular and linear Ag trimers with total charges of +1 and –1, respectively. These systems define an ill-posed learning problem for models which are unaware of charge in a similar way to the toy datasets from Section 3.1.
Sodium chloride clusters, $\mathbf{Na_{8/9}Cl_{8}^{+}}$. This benchmark contains a positively charged $\mathrm{Na_{9}Cl_{8}^{+}}$ cluster and the system resulting from the removal of a neutral sodium atom, $\mathrm{Na_{8}Cl_{8}^{+}}$. The +1 total charge of the cluster is maintained, resulting in a redistribution of charge when the sodium atom is removed.
In line with previous work, we trained extended TensorNet on 90% data points, being the remaining ones used as a separate validation set. Comparative results for 4D-HDNNP21, PANNA-LR22, CACE-LR 28, SpookyNet16 and TensorNet can be found in Table 3. In the case of 4G-HDNNP, PANNA-LR ($\gamma_{q}>0$) and SpookyNet, the models are trained to reproduce DFT atomic reference charges and include an explicit electrostatic term in the energy prediction that makes use of these. In SpookyNet, nuclear repulsion and D4 dispersion terms are also included. We emphasize that, in contrast, TensorNet was directly trained on energy and force labels without making use of any charge predictions or prior physical terms. CACE-LR makes use of Latent Ewald Summation 27, which learns hidden ’latent charges’ from atomic features without explicit reference to any specific charge definition, and uses these to compute long-range interactions via Ewald summation while avoiding the need for charge equilibration. However, one can consider the assumed interaction in reciprocal space as an electrostatic-like ansatz. Therefore, we compared our approach to a wide diversity of methods in the literature.
Our model performs competitively or outperforms SpookyNet, the current best performing architecture. The baseline TensorNet already shows strong performance on the $\mathrm{C_{10}H_{2}/C_{10}H_{3}^{+}}$ and $\mathrm{Na_{8/9}Cl_{8}^{+}}$ systems even without charge awareness. This is because these systems can be distinguished structurally: $\mathrm{C_{10}H_{3}^{+}}$ has an additional proton compared to $\mathrm{C_{10}H_{2}}$, while $\mathrm{Na_{9}Cl_{8}^{+}}$ has one more sodium atom than $\mathrm{Na_{8}Cl_{8}^{+}}$. However, for $\mathrm{Ag_{3}^{+/-}}$, where three silver atoms form either a cation or anion, the baseline model fails to distinguish between charge states, leading to large errors (1683.11 meV/atom). The inclusion of charge awareness maintains the strong performance on structurally distinguishable systems while dramatically improving predictions for $\mathrm{Ag_{3}^{+/-}}$ (0.888 meV/atom), demonstrating its ability to resolve input degeneracy when geometric information alone is insufficient.
Furthermore, we compared the speed performance of SpookyNet and TensorNet for a carbon chain conformation, using the ASE Calculator implementation 34, 35 for SpookyNet and the TorchMD-Net external calculator for TensorNet. The speed benchmark was conducted on an NVIDIA RTX 4090, with both models calculating energy and forces over 100 runs. TensorNet achieved a mean runtime of 1.462 ms per inference step, while SpookyNet recorded 1.751 ms. Notably, while SpookyNet’s long-range electrostatics scale as $O(N^{2})$, our method scales as $O(N)$.
Table 3: Charged benchmark systems $\mathbf{C_{10}H_{2}/C_{10}H_{3}^{+}}$, $\mathbf{Ag_{3}^{+/-}}$ and $\mathbf{Na_{8/9}Cl_{8}^{+}}$. Energy (E) and forces (F) root mean squared errors in meV/atom and meV/Å, with standard deviation over 3 splits between parentheses for TensorNet, on the validation set consisting of 10% of data points. $Q$ refers to total charge, with factor corresponding to weights $\lambda_{Q}=\tilde{\lambda}_{Q}=0.1$. Results for 4G-HDNNP and SpookyNet have been taken from  16, PANNA-LR results have been extracted from  22 corresponding to the best-performing model trained on partial charges, $\gamma_{q}>0$ as denoted in the reference, while CACE-LR results have been obtained from  28.
Model $\mathbf{C_{10}H_{2}/C_{10}H_{3}^{+}}$ $\mathbf{Ag_{3}^{+/-}}$ $\mathbf{Na_{8/9}Cl_{8}^{+}}$
4G-HDNNP [21] E (meV/atom) 1.194 1.323 0.481
F (meV/Å) 78.00 31.69 32.78
PANNA-LR [22] E (meV/atom) 1.17 0.80 0.40
F (meV/Å) 79 20 19
CACE-LR [28] E (meV/atom) 0.73 0.162 0.21
F (meV/Å) 36.9 29.0 9.78
SpookyNet [16] E (meV/atom) 0.364 0.220 0.135
F (meV/Å) 5.802 26.64 1.052
TensorNet E (meV/atom) 0.124 (0.001) 1683.11 (203.36) 0.103 (0.003)
F (meV/Å) 1.88 (0.02) 912.67 (39.20) 1.73 (0.06)
TensorNet+$0.1Q$ E (meV/atom) 0.124 (0.001) 0.888 (0.019) 0.102 (0.003)
F (meV/Å) 1.90 (0.15) 20.70 (0.77) 1.78 (0.09)
3.4 Solvated protein fragments
The solvated protein fragments dataset, introduced in Ref 5, includes structures for all possible hydrogen-saturated covalently bonded fragments with up to eight heavy atoms (C, N, O, S) derived from proteins. It accounts for different charge states of aminoacids due to (de)protonation, with total charges up to ±2. The dataset features solvated variants with varying water molecule numbers, and randomly sampled dimer interactions of protein fragments, along with pure water structures with up to 40 molecules. Multiple conformations are included, resulting in more than 2.7M structures with reference energies, forces, and dipole moments.
In line with Ref  5, we trained TensorNet with the inclusion of total charge ($\lambda_{Q}=\tilde{\lambda}_{Q}=0.1$) on the energy and forces, excluding dipoles, of 2560k structures, with 100k of these being used for validation, being the remaining ones used for testing. Results are found in Table 4, and demonstrate that the modified TensorNet significantly outperforms the more accurate ensemble of five PhysNet models, which use explicit electrostatic and dispersion energy terms in the predictions, particularly in energies, nearly halving its error.
Table 4: Solvated protein fragments results. Comparison of mean absolute errors on the solvated protein fragments test set for energies (E) and forces (F). $Q$ refers to total charge, with the factor corresponding to $\lambda_{Q}=\tilde{\lambda}_{Q}=0.1$. PhysNet-ens5 denotes an ensemble of five PhysNet models. Results have been taken from  5.
PhysNet PhysNet-ens5 TensorNet+$0.1Q$
E 1.03 kcal/mol 0.95 kcal/mol 0.55 (0.04) kcal/mol
F 0.88 kcal/mol/Å 0.72 kcal/mol/Å 0.63 (0.09) kcal/mol/Å